A CHIRPED PULSE FTMW STUDY OF THE STRUCTURE OF PHENOL DIMER

\underline{AMANDA L. STEBER}, JUSTIN L. NEILL, DANIEL P. ZALESKI, and BROOKS H. PATE, Department of Chemistry, University of Virginia, Charlottesville, VA 22904; ALBERTO LESARRI, Departamento Qumica Fsica y Qumica Inorgnica,Facultad de Ciencias, Universidad de Valladolid, 47011 Valladolid, Spain.

Phenol dimer has been studied extensively and is considered a benchmark molecular complex for ab initio theory due to a long range dispersion interaction between the rings as well as an intermolecular hydrogen bond. Previously, the structure had been determined using RCS^{a}^{,}^{b} and high resolution UV measurements;^{c} however, several assumptions were integrated into the structure because a full isotopically substituted structure could not be determined. In this study, the rotational spectrum of the dimer as well as ^{13}C and ^{18}O isotopologue spectra that were seen in natural abundance were obtained using chirped pulse Fourier transform microwave spectroscopy (CP-FTMW). The structure was determined using both linear least squares fitting (r_{0} structure) and the Kraitchman substitution analysis (r_{s} structure). Ab initio calculations were performed for the dimer using MP2/cc-pVTZ cp^{d}, B3LYP/6-31G(d,p), M06-2X/6-31G(d,p), and M06-2X/6-311++G(d,p), while CCSD calculations are currently under way. Changing the level of theory and the basis set dramatically changes the structure. The MP2 calculation underestimates the hinge angle (C-O-O-C dihedral angle), while the B3LYP overestimates it. The M06-2X calculations seem to give the best cost-to-benefit ratio when compared to the r_{s} structure, but they show poorer agreement with increasing basis set size.

aL. L. Connell, S. M. Ohline, P. W. Joireman, T. C. Corcoran, and P. M. Felker, J.Chem.Phys. 96, 2585 (1992)

bA. Weichert, C. Riehn, and B. Brutschy, J. Phys. Chem. A 105, 5679 (2001)

cM. Schmitt, M. Böhm, C. Ratzer, D. Krügler, K. Kleinermanns, I. Kalkman, G. Berden, and W. L. Meerts, Chem. Phys. Chem.

7, 1241 (2006)

dP. Jurecka, J. Sponer, J. Cerny, and P. Hobza, Phys. Chem. Chem. Phys. 8, 1985 (2006)